a car moving at a speed of 25 m/s in the postive direction decelerates to a stop in 5.0 s. The average acceleration of the car over the 5.0s is
1
Expert's answer
2020-11-18T09:24:03-0500

Vi = 25 m/s

Vf = 0 m/s

t = 5.0 s

Average acceleration:

"a = \frac{V_f - V_i}{t}"

"a = \frac{0 - 25}{5} = \frac{-25}{5} = - 5 \;m/s^2" (negative sign indicates deceleration)